• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name silwal

    Log entry time 03:21:34 on November 8,2009

    Entry number 298252

    keyword=target moved to empty

    target moved to empty because we dont have the spectrometers right now. Left arm down for the night, right arm dipole still down, might be back up later on.

    FIGURE 1